a) What is the magnitude of the John's acceleration?

b) What is the direction of the John's acceleration?

c) If the pulling force is applied constantly, where will Jane and John meet? from John
Answer #1

The answer is simple.

If Jane's acceleration is 0.8 then the force is 54*0.8 = 43.2 N

John experiences the equal and opposit force so his acceleration is 43.2/63 = 0.685 in the direction opposite to Jane's acceleration.

Since the net force on the system is 0, they will meet at the system's centre of mass which is (from John's original position) = (63*0 + 54*10)/(54+63) =4.61 meters from John